OpenStudy (anonymous):

Solve the equation for x. Round your answer to two decimal places. Do not include "x =" in your answer. 4*e^x = 17

OpenStudy (valpey):

\[4e^x=17\] \[e^x=\frac{17}{4}\] \[\ln{e^x}=\ln{\frac{17}{4}}\] \[x=\ln{\frac{17}{4}}\] Use a calculator to find x = ....